OpenStudy (anonymous):

Write an equation in standard form for the ellipse with center (0, 0), vertex (9, 0) and co-vertex (0, 5).

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you got the center at \((0,0)\) so you know it looks like \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] what you are missing is \(a, b\)

OpenStudy (anonymous):

so what should a and b be??

OpenStudy (anonymous):

vertex is \((9,0)\) so i guess the other one is \((-9,0)\) making \(a=9\)

OpenStudy (anonymous):

that isn't one of my options :/

OpenStudy (anonymous):

and then of course \(b=5\) making your equation \[\frac{x^2}{81}+\frac{y^2}{25}=1\]

OpenStudy (anonymous):

i think it is right, we can check it if you like

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=ellipse+x^2%2F81%2By^2%2F25%3D1 looks good to me

OpenStudy (anonymous):

a. \[\frac{ x^2 }{ 49 }+\frac{ y^2 }{ 36 }=1\] b. \[\frac{ x^2 }{ 49 }+\frac{ y^2 }{ 13 }=1\]

OpenStudy (anonymous):

only those two? i still like my answer

OpenStudy (anonymous):

check the wolfram link see that the vetices are as advertised

OpenStudy (anonymous):

the other two are written wrong like y before x and just do not look right

OpenStudy (anonymous):

sorry but i still like my answer it has vertices \((9,0)\) and \((-9,0)\) covertices \((0,5)\) and \((0,-5)\) and center \((0,0)\) maybe there is a typo somewhere in the problem

OpenStudy (anonymous):

\[\frac{x^2}{49}+\frac{y^2}{36}=1\] has vertices \((7,0)\) and \((-7,0)\) covertices \((0,6),(0,-6)\)

OpenStudy (anonymous):

okay i will go with that one thank you so much for your help. I will be emailing my teacher ASAP lol!!